Suppose you analyzed the average monthly credit card bill of Visa credit card customers using two randomly selected samples: one with 144 Visa credit card customers (Sample #1) and one with 100 Visa credit card customers (Sample #2). Every credit card customer in Sample #1 had a monthly credit card bill of $1,812. Sample #2, on the other hand, had an average monthly credit card bill of $1,868 with a variance of 1,302,388 ($)^2. If you combined these samples into one... Question 2What would be the combined monthly credit card bill average (in dollars)? Question 3AWhat is the combined monthly credit card bill median (in dollars)? Question 3BWhat is the shape of the combined monthly credit card bill distribution (in dollars)? Question 4 What would be the combined monthly credit card bill standard deviation (in dollars)?
Answer #1

Solution2:

combined monthly credit card bill average=n1x1bar+n2x2bar/n1+n2

x1bar=1812

x2bar=1868

n1=144

n2=100

combined monthly credit card bill average=144*1812+100*1868/144+100

=1834.95

Solution3a:

large samples follows normal distribution

mean=median

combined monthly credit card bill median=1834.95

Solution3b;

combined sample =144+100=244

244>30

large sample follows normal distribution

shape :symmetrical

Solution4:

combined monthly credit card bill standard deviation=sqrt(s1^2+s2^2/2

=sqrt(1302388/2)

=806.97